OpenStudy (anonymous):

Determine if V is a vector space when: x+y=xy cx=x^c If not state all the vector axioms it fails

OpenStudy (anonymous):

If yes verify each vector space axiom

OpenStudy (turingtest):

here's at least one that fails\[c(\vec u+\vec v)\neq c\vec u+c\vec v\]

OpenStudy (anonymous):

ok what abt tat scaler multiplication?

OpenStudy (phi):

x^c is not a linear operation

OpenStudy (anonymous):

okkkkk

OpenStudy (turingtest):

actually it's consistent though... \[c(\vec u+\vec v)= c\vec u+c\vec v\]because we have\[c(\vec u+\vec v)=(uv)^c=u^cv^c= c\vec u+c\vec v\]so I don't see what's wrong with it

OpenStudy (anonymous):

well it is definitely not closed under scaler multilication though

OpenStudy (turingtest):

how so?

OpenStudy (anonymous):

because x^c is exponential

OpenStudy (turingtest):

but that doesn't necessarily take it out of V

OpenStudy (turingtest):

*the vector space I mean

OpenStudy (anonymous):

oh ya? ok I dont get a thing, I feel so stupid

OpenStudy (turingtest):

aha! it is a vector space I thought I'd seen it befor check it out, the zero vector turns out to be 1 I think http://tutorial.math.lamar.edu/Classes/LinAlg/VectorSpaces.aspx

OpenStudy (anonymous):

yay lol :D

OpenStudy (phi):

@Tur how do you get from (uv)^c to cu + cv?

OpenStudy (turingtest):

check out example 5

OpenStudy (anonymous):

yup i see it. Thanks :D

OpenStudy (turingtest):

@ phi that is what you get when you put the two rules together the exponent is distributed because vector addition is defined as multiplication

OpenStudy (phi):

nvm, i'll look at your link...

OpenStudy (anonymous):

LOL i htink i am fundamentally lacking basic knowledge

OpenStudy (turingtest):

I'm not so good at this stuff, I had it wrong at first as well